# Query-Focused Retrieval Heads Improve Long-Context Reasoning and Re-ranking

### v1

@@ Abstract
Wuwei Zhang  Fangcong Yin  Howard Yen  Danqi Chen  Xi Ye Princeton Language and Intelligence, Princton University The University of Texas at Austin {wuwei.zhang,hyen,danqic}@cs.princeton.eduxi.ye@princeton.edu fangcongyin@utexas.edu
@@ 2 Background: Retrieval Heads
$\vspace{-0.5em}\mathrm{Retrieval\_Score}(h)=\frac{|g_{h}\cap a|}{|a|},$ (1)
@@ 4.2 Applications
Long-context language models often struggle with performance degradation when processing long context Yen et al. (2025); Ye et al. (2025); Liu et al. (2024). To address this, we integrate QRRetriever within a retrieval-augmented generation (RAG) framework. Given a long-context input and a query, we segment the input into smaller chunks and use QRRetriever to score and subsequently extract the most relevant ones. The extracted context are concatenated to create a reduced context, that is then given to the LM for generating the final answer in another forward pass.
@@ 5.2 Long-Context Multi-Hop Reasoning
As shown in Table 1, on CLIPPER, Llama-3.2-3B-Instruct achieves a Recall@10 of 86.1, closely matching the 88.3 score of the much larger LlamA-3.1-70B. However, Llama-3.2-3B only achieves a final end-to-end performance of 24.0, largely lagging 70B’s performance of 76.7. We hypothesize that the long-context limitations of compact models stem more from their generation capabilities than from their retrieval abilities. These findings open up promising future directions. Compact LMs could serve as efficient long-context retrievers, paired with larger models for the actual generation.
@@ 5.3 Passage Re-Ranking
For BEIR, we utilize the 128 (held-out) data points from NQ and use them for on all other domains zero-shot.
@@ Results.
Table 2 summarizes the BEIR results, demonstrating the strong effectiveness of QRRetriever as a general-purpose retriever. For models under 10B parameters, QRRetriever consistently outperforms other baselines. With Llama-3.1-8B, it achieves an average score of 49.7, outperforming RankGPT by 3.4 points and ICR by 1.2 points. For the larger Llama-3.1-70B model, QRRetriever significantly surpasses ICR, though it generally lags RankGPTBubble (which require over 200 generation calls). Nevertheless, QRRetriever achieves the best performance on several domains, such as SciFact and FEVER. In addition, we find that QRRetriever directly benefits from the general-purpose capabilities of strong LLMs, outperforming popular dual encoders and cross-encoders that fine-tune various base models to improve retrieval performance.
@@
Llama-8B Random Heads 37.5 59.8
Full Heads 42.8 73.2
@@
QRHead 47.5 85.6
Qwen-7B Random Heads 19.9 57.2
@@ 7 Related Work
Interpretability studies have shown that many core behaviors of LMs, including in-context learning Olsson et al. (2022); Todd et al. (2024); McDougall et al. (2023) and retrieval Wu et al. (2025b), can be traced to specialized transformer modules Meng et al. (2022); Dai et al. (2022); Stolfo et al. (2024). Techniques have been proposed to localize such modules with a small amount of data Meng et al. (2022); Geiger et al. (2024); Bhaskar et al. (2024), and to intervene on them for control Li et al. (2023); Yin et al. (2024); Huang et al. (2025) or efficiency Tang et al. (2025); Xiao et al. (2025). However, only a few works Zhao et al. (2025) have examined attention head specialization in long-context settings, where attention is known to be not robust Liu et al. (2024); Xiao et al. (2024a), and it is an open question if intervening the localized modules is crucial in practical settings Hase et al. (2023); Wang and Veitch (2024). Our work contributes to this line of research by finding better specialized set of attention heads that explain the model behavior for query-focused long-context retrieval, and that can be practically useful for zero-shot efficient retrieval.
@@ Appendix A Details about Evaluation Datasets
targets narrative claim verification—a challenging long-context reasoning task that requires verifying claims over entire books, with an average length of $90$K tokens and $23$ chapters. In CLIPPER, data is split at the chapter level, with each chapter treated as an individual document during retrieval.
@@ Appendix C Details about Dual Encoder and Cross Encoder Baselines on BEIR
Appendix D License of Datasets
@@
Appendix E Computational Resources and Model Sizes
@@
Appendix F Potential Risks of Our Work

### v2

@@ Abstract
Wuwei Zhang   Fangcong Yin   Howard Yen   Danqi Chen   Xi Ye Princeton Language and Intelligence, Princeton University The University of Texas at Austin {wuwei.zhang,hyen,danqic}@cs.princeton.edu  xi.ye@princeton.edu fangcongyin@utexas.edu
@@ 2 Background: Retrieval Heads
$\vskip-5.0pt\mathrm{Retrieval\_Score}(h)=\frac{|g_{h}\cap a|}{|a|},$ (1)
@@ 4.2 Applications
Long-context language models often struggle with performance degradation when processing long context Yen et al. (2025); Ye et al. (2025); Liu et al. (2024a). To address this, we integrate QRRetriever within a retrieval-augmented generation (RAG) framework. Given a long-context input and a query, we segment the input into smaller chunks and use QRRetriever to score and subsequently extract the most relevant ones. The extracted context are concatenated to create a reduced context, that is then given to the LM for generating the final answer in another forward pass.
@@ 5.2 Long-Context Multi-Hop Reasoning
As shown in Table 1, on LongMemEval, Llama-3.2-3B-Instruct achieves a Recall@10 of 86.6, closely matching the 88.5 score of the much larger LlamA-3.1-70B. However, Llama-3.2-3B only achieves a final end-to-end performance of 47.7, largely lagging 70B’s performance of 67.7. We hypothesize that the long-context limitations of compact models stem more from their generation capabilities than from their retrieval abilities. These findings open up promising future directions. Compact LMs could serve as efficient long-context retrievers, paired with larger models for the actual generation.
@@ 5.3 Passage Re-Ranking
For BEIR, we utilize the 256 (held-out) data points from NQ and use them for on all other domains zero-shot.
@@ Results.
Table 2 summarizes the BEIR results, demonstrating the strong effectiveness of QRRetriever as a general-purpose retriever. For models under 10B parameters, QRRetriever consistently outperforms other baselines. With Llama-3.1-8B, it achieves an average score of 50.0, outperforming RankGPT by 3.7 points and ICR by 1.5 points. For the larger Llama-3.1-70B model, QRRetriever significantly surpasses ICR, though it generally lags RankGPTBubble (which require over 200 generation calls). Nevertheless, QRRetriever achieves the best performance on several domains, such as SciFact and FiQA. In addition, we find that QRRetriever directly benefits from the general-purpose capabilities of strong LLMs, outperforming popular dual encoders and cross-encoders that fine-tune various base models to improve retrieval performance.
@@
Llama-8B
Random Heads 37.5 59.8
Full Heads 42.8 77.0
@@
QRHead 47.5 85.5
Qwen-7B
Random Heads 19.9 57.2
@@ 6 Analysis
In this section, we analyze the key advantages of QRHead (e.g., length generalization) and examine the factors underlying its effectiveness. Additional analyses on the impact of varying the number of heads (Appendix F) and inference latency (Appendix E) are provided in the appendix.
@@ 7 Related Work
Interpretability studies have shown that many core behaviors of LMs, including in-context learning Olsson et al. (2022); Todd et al. (2024); McDougall et al. (2023) and retrieval Wu et al. (2025b), can be traced to specialized transformer modules Meng et al. (2022); Dai et al. (2022); Stolfo et al. (2024). Techniques have been proposed to localize such modules with a small amount of data Meng et al. (2022); Geiger et al. (2024); Bhaskar et al. (2024), and to intervene on them for control Li et al. (2023); Yin et al. (2024); Huang et al. (2025) or efficiency Tang et al. (2025); Xiao et al. (2025); Liu et al. (2024b). However, only a few works Zhao et al. (2025) have examined attention head specialization in long-context settings, where attention is known to be not robust Liu et al. (2024a); Xiao et al. (2024a), and it is an open question if intervening the localized modules is crucial in practical settings Hase et al. (2023); Wang and Veitch (2024). Our work contributes to this line of research by finding better specialized set of attention heads that explain the model behavior for query-focused long-context retrieval, and that can be practically useful for zero-shot efficient retrieval.
@@ Appendix A Details about Evaluation Datasets
targets narrative claim verification—a challenging long-context reasoning task that requires verifying claims over entire books, with an average length of 9090K tokens and $23$ chapters. In CLIPPER, data is split at the chapter level, with each chapter treated as an individual document during retrieval.
@@ Appendix C Details about Dual Encoder and Cross Encoder Baselines on BEIR
Appendix D Prompt Templates
We provide prompt templates used in our experiments for ICR and QRRetriever in Figure 5 and rankGPT in Figure 6.
{prompt_prefix} Here are some paragraphs:
[1] {Title 1 (if available)}
{Paragraph text 1}
[2] {Title 2 (if available)}
{Paragraph text 2}
Please find information that is relevant to the following query in the paragraphs above.
Query: {query}{prompt_suffix}
Figure 5: Prompt used for ICR and QRRetriever.
{prompt_prefix} This is an intelligent assistant that can rank passages based on their relevancy to the query.
The following are {N} passages, each indicated by a numbered identifier [i]. I can rank them based on their relevance to the query: "{query}"
[1] {Title 1 (if available)}
{Paragraph text 1}
[2] {Title 2 (if available)}
{Paragraph text 2}
The search query is: "{query}". I will rank the {N} passages above based on their relevance to the search query. The passages will be listed in descending order using identifiers, the most relevant passages should be listed first and the output format should be [] > [] > etc, e.g., [1] > [2] > etc. Be sure to list all {N} ranked passages and do not explain your ranking until after the list is done. {prompt_suffix} Ranked Passages: [
Figure 6: Prompt used for rankGPT.
Appendix E Inference Time Comparison between QRRetriever, ICR, and RankGPT
We compare the latency of QRRetriever with other LLM-based re-rankers (ICR and RankGPT). In Table 7, we report both latency and performance of Llama-3.1-8B and Llama-3.1-70B on NQ dataset. Compared to RankGPT, QRRetriever is significantly more time-efficient. This is primarily because (1) QRRetriever avoids autoregressive generation, and (2) it does not rely on bubble sort, which requires multiple rounds of generation to compare elements sequentially within each bubble.
Avg Latency (s/sample) NQ
nDCG@10
Model: LLama-3.1-8B-Inst
RankGPTBubble 14.5 53.7
ICR 2.2 54.0
QRRetriever 2.2 58.6
Model: Llama-3.1-70B-Inst
RankGPTBubble 63.9 58.4
ICR 13.7 57.9
QRRetriever 13.7 62.1
Table 7: Inference time and retrieval performance comparison on NQ.
Appendix F Determine the Number of Heads Used
We select the number of heads based on the sparsity level reported in prior work on retrieval heads Wu et al. (2025b). In Table 8, we include additional results on BEIR NQ using varying numbers of heads. The results show that performance remains consistently strong when using a small, top-ranked subset of heads, but degrades as more heads are included. This trend aligns with the sparsity property observed in retrieval heads.
Model / #Heads NQ (nDCG@10)
Llama-3.1-8B (#heads = 1024)
16 ($\sim$1.5%) 58.6
64 ($\sim$6%) 57.4
256 ($\sim$20%) 56.6
Llama-3.1-70B (#heads = 5120)
32 ($\sim$0.5%) 62.1
128 ($\sim$2%) 61.9
512 ($\sim$10%) 61.0
Table 8: Performance with different numbers of selected heads.
Appendix G License of Datasets
@@
Appendix H Computational Resources and Model Sizes
@@
Appendix I Potential Risks of Our Work